# A mathematical theory for understanding when abstract representations emerge in neural networks

### v1

@@ II.4 The optimality condition
To compute the $PS$ [Eq. (7)] corresponding to the ground state of effective system, we want to find the representation kernel $K[\rho_{M}^{*}]$ corresponding to the global minima $\rho_{M}^{*}$ of Eq. (14), with the constraint that $\rho_{M}^{*}$ is a sum of $M$ Dirac-delta measures and $\text{supp}\rho_{M}^{*}\subseteq\text{Range}K_{X}$. This optimization is challenging to perform directly, as the space of finite $M$-sum Dirac delta measures is not convex. To address this, we relax the constraint by enlarging the optimization domain to include all finite positive measures supported on $\text{Range}K_{X}\subseteq\mathbb{R}^{P}$, which also includes continuous measures. We denote this new space of measures  [^2] To be technically correct, we also require all 2nd-order moments of $\rho$ are finite in $M_{+}(X)$. This would ensure all the integrals in $E[\rho]$ are finite., as $M_{+}(K_{X})$. By definition, $M_{+}(K_{X})$ is convex.
@@
This is a convex functional on $M_{+}(X)$: since $M_{+}(X)$ is a convex set (equipped with the usual sum between measures), $E[\rho]$ is a convex function on $M_{+}(X)$ if and only if $\forall\rho_{1},\rho_{2}\in M_{+}(X),\,f(t)\equiv E[t\rho_{1}+(1-t)\rho_{2}]$ is a convex function for $t\in[0,1]$ [12]. The latter can be checked by computing the 2nd derivative of $f(t)$ (SI §1.3),
@@
= $"="\text{ holds}\,$ $\Rightarrow\,\mathbf{\forall\,h}\in\text{supp}(\rho_{*}),$ (17)
where $K[\rho_{*}]=\int\phi(\mathbf{h})\phi(\mathbf{h})^{T}\,d\rho_{*}(\mathbf{h})$ is the representation kernel for $\rho_{*}$. Because of convexity, any solution $\rho_{*}$ satisfying the KKT condition is automatically a global minimum of Eq. (15).
This KKT condition shows that any global minimum $\rho_{*}$ of Eq. (15) must satisfy the inequality in Eq. (II.4) and its support is confined to those vectors $\mathbf{h}$ for which the equal sign is attained. The two conditions can be interpreted as a mean-field description [69] of the effective $M$-neuron system (Fig. 3F). To see this, we define the single-neuron mean-field energy,
@@
And the KKT conditions [Eq. (II.4)] are equivalent to
@@
$\text{supp}(\rho_{*})\subseteq\underset{\mathbf{h}\in\text{Range}K_{X}}{\text{argmin}}E(\mathbf{h};\rho_{*}),$ (19)
for any optimal solution $\rho_{*}$. The support of $\rho_{*}$ contains the optimal preactivation patterns (or single-neuron tuning) for training data, which we denote as $A(K_{X},K_{Y})$.
@@
These mean-field equations [Eq. (II.4)] need to be solved in a self-consistent manner [69, 47, 46, 63] (See SI §1.4). For arbitrary input and output data $K_{X},K_{Y}$, and nonlinear activation function $\phi$, these equations are systems of nonlinear equations that need to be solved numerically. Fortunately, for training data related abstract representation (Section II.1), the mean-field equations [Eq. (II.4)] can be solved exactly as we will present below.
As a final note, although the above optimization is performed over all positive measures in $M_{+}(X)$ and in general, some of these measures cannot be attained by a finite-size system, in many cases, the global minimizer $\rho_{*}$ is a finite sum of Dirac-delta measures. As we show in the next section, this fact allows us to study the optimal neural representations in a finite-width network by optimizing Eq. (15) over the infinite-dimensional space $M_{+}(X)$. And the latter problem is a convex problem [Eq. (14)].
@@ III Whitened and target-aligned inputs lead to abstract representation in ReLU network
For these input geometries, we find that all the solutions of the mean-field problem [Eq. (II.4)] correspond to abstract hidden representations. We summarize the key steps of the argument here, with detailed derivations provided in SI §2. We start with the scenario where each binary class contains a single data point $(n=1)$ and then extend the results to multi-element classes $(n\geq 2)$.
@@ III.1 Whitened input + single-element class ($n=1$)
Both matrices are positive definite. So the constraint on $\mathbf{h}$ in Eq. (II.4) is trivial, $\text{Range}K_{X}=\mathbb{R}^{P}$.
@@
where $\mathbf{h}_{+}=[\mathbf{h}]_{+}$ and $\mathbf{h}_{-}=[-\mathbf{h}]_{+}$ are the positive and negative components of the vector $\mathbf{h}$. Using this inequality, we can find the minimum of $E(\mathbf{h};\rho)$ [as required in Eq.(II.4)] via minimizing $E_{r}(\mathbf{h}_{+};\rho)$ over all the nonnegative vectors $\mathbf{h}_{+}\geq 0$.
@@
Here $\lambda_{1,2}$ are assumed to be small to ensure $b_{*}>0$, i.e. $\lambda_{1}\lambda_{2}<\frac{P+2}{P(P+1)}$. Interestingly, the solutions of the mean-field equation $\rho_{*}$ [Eq.(II.4)], or equivalently the global minimizers of the loss [Eq.(15)], are not unique. But they all correspond to the identical representation kernel given by Eq.(21).
Figure 4: Task-optimized ReLU network exhibits abstract representation for whitened and target-aligned inputs. (A) The training loss and the parallelism score of the hidden representation are plotted against the number of training epochs for the whitened input. The training is through a gradient descent algorithm. After training, the network performs the task perfectly with zero training loss and has an abstract hidden representation ($PS\rightarrow 1$). (B) The optimal hidden kernel predicted by theory ($K_{theory}$ given by Eq. (21)) is aligned with the one found in numerical simulation ($K_{sim}$) for different output dimensions $d_{Y}$. (C) The parallelism score of the hidden representation after training as a function of the input-output alignment. See SI §5 for the definition of input-output alignment. Each point in the plot represents a specific input geometry with a randomly sampled input kernel. The red dot indicates the point for the whitened input kernel. For inputs that are more aligned to the output than the whitened one, the PSPS is close to $1$. (D-E) Modularity of the single-neuron tuning in the hidden layer is captured by the preactivation vector $\mathbf{h}$ and weight vector $\mathbf{v}$ to the output layer for each hidden neuron. Here $d_{Y}=2$. (E) Each hidden neuron only has nonzero output weights to a single output unit, suggesting that different neurons in the hidden layer are used to ”read out” different output labels. (F) The first three principal components of the neural preactivation space (the same space as Fig. 3B). The preactivation vectors of the hidden neurons concentrate along $P=4$ distinct directions, as predicted by Eq. (23).
The support of $\rho_{*}$ gives the set of optimal preactivation vectors in the hidden layer $A(K_{X},K_{Y})$, which is found to consist of $2d_{Y}$ line rays in the nonnegative orthorant $\mathbb{R}_{\geq 0}^{P}$ (SI §2.2)
@@
As half of the components in $\mathbf{v}_{i}$ are +1+1 and the other ones are 1-1, these optimal preactivation vectors have exactly half of the components to be $2$ and the other ones are zero.
@@
Altogether, when $M\geq 2d_{Y}$ and the input is whitened, we find the optimal hidden representation as in Eq.(21) and that the neurons in the hidden layer clusters into $2d_{Y}$ groups given by Eq.(23).
@@ III.2 Whitened input + multi-element class ($n\geq 2$)
In SI §2.3, we solve the mean-field problem [Eq. (II.4)] and find the optimal representation kernel is indeed similar as before
@@
This optimal representation kernel [Eq. (25)] has an intriguing geometric interpretation: the between-class correlation matrix $b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y}^{0})$ is the same as for the single-element class case [Eq. (21)] and the within-class correlation matrix is given by $\mathbf{1}\mathbf{1}^{T}$. This form of within-class correlation means the neural representation of all data points from the same binary class “collapse” into a single point in the hidden layer, as also indicated in the set of optimal preactivations [Eq. (27)]. Such a property closely resembles the neural collapse phenomena reported in previous studies [72, 101, 92, 89, 88, 37]. In SI §5.1, we discuss the similarity and difference between our results and the work related to neural collapse. For ReLU nonlinearity, this ”within-class collapse” property can be directly proved via Jensen’s inequality (SI §2.4). The fact that the optimal preactivation patterns in the hidden layer clusters into only a few directions [Eq. (23) and Eq. (25)] is also known as the quantization phenomenon for ReLU networks [55].
@@ III.3 Target-aligned input
Unlike the fully orthogonalized input before, the input here has positive (homogeneous) correlations if $c_{0}$ is large. Interestingly, we find (SI §2.6) that this introduced correlation does not change the form of the optimal kernel $K[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y})$, and only modifies the coefficient $b_{*}$
@@
The optimal kernel in this case has the same form as for orthogonalized inputs [Eq. (25)], $K[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y}^{0})\otimes\mathbf{1}\mathbf{1}^{T}$, but with an optimal coefficient (SI §2.6)
@@
Altogether, this shows that for target-aligned inputs (Eq.(29)), the optimal representation kernel [Eq. (28)] does not depend on the magnitude of orthogonal components $c_{j}$ for all $j=d_{Y}+1,..,P-1$. However, this property does not hold if the target-aligned condition [Eq.(29)] is not satisfied, i.e. when the orthogonal component is large, $c_{j}>c_{Y}$ (Fig.5C). The underlying intuition is that for nonlinear networks, the output-aligned component and the output-orthogonal component in the input ”compete” with each other to form the hidden representation. In contrast, a linear network only uses the output-aligned component to form the hidden representation.
@@ III.4 $PS$, single-neuron tuning and weight matrices for the optimal solution
From the optimal representation kernel, we can compute its $PS$, denoted $PS(K[\rho_{*}])$. By translation-and scale-invariance of $PS$ (Section II.2), $PS(K[\rho_{*}])=PS(K_{Y})$. By the definition of $K_{Y}$, the prototype representation [Eq. (6)] of each binary class corresponds to each vertex of a $d_{Y}$-dimensional hypercube (Fig. 2). So the representation change for the $k$th binary latent label is aligned with the $k$th axis of the hypercube (Fig.4), independent of other latent labels: $PS_{k}(\alpha_{1},\alpha_{2})\equiv\frac{\Delta\mathbf{r}(k;\alpha_{1})\cdot\Delta\mathbf{r}(k;\alpha_{2})}{\left\|\Delta\mathbf{r}(k;\alpha_{1})\right\|_{2}\left\|\Delta\mathbf{r}(k;\alpha_{2})\right\|_{2}}=1$. Therefore, the overall parallelism score $PS(K_{Y})=1=PS(K[\rho_{*}])$. And the optimal representation kernel [Eq. (24) and (31)] corresponds to an abstract representation.
Next, we look at the responses of individual neurons in the optimal solution $\rho_{*}$. From Eq. (23) and Eq. (27), the optimal preactivation of each neuron has the form $\mathbf{h}=\alpha(1\pm\mathbf{v}^{0}_{i})\otimes\mathbf{1}$ for some $i\in\{1,2,3..,d_{Y}\}$ and $\alpha>0$ (we focus on neurons having nonzero preactivation patterns). The $P$ components of $\mathbf{h}\in\mathbb{R}^{P}\simeq\mathbb{R}^{2^{d_{Y}}}\otimes\mathbb{R}^{n}$ are this neuron’s preactivations for the $P$ training data points [Eq.(11)]. Based on the definition of $\mathbf{v}^{0}_{i}$ (Section II.1 and SI §1.1), the neuron with $\mathbf{h}=\alpha(1+\mathbf{v}^{0}_{i})\otimes\mathbf{1}$ (or $\mathbf{h}=\alpha(1-\mathbf{v}^{0}_{i})\otimes\mathbf{1}$) has nonzero preactivations exactly for those training data points whose $i$th output label are positive (or negative). So all the neurons in the hidden layer are divided into $2d_{Y}$ groups: neurons in each group respond only based on a single output label (Fig.5D-F). Note that a random network would not have a modular response property like this, and such modularity is a consequence of task training.
Since the preactivation of every neuron is a linear combination of $\mathbf{1\otimes\mathbf{1}}$ and $\mathbf{v}_{i}\otimes\mathbf{1}$ for some $k$, the optimal preactivation matrix $H_{*}$ [Eq. (11)] has the following form
@@
where $\mathbf{A}\in\mathbb{R}^{(d_{Y}+1)\times M}$ is the matrix encoding the coefficients $\alpha$ for every neuron. Importantly, $H_{*}\in\mathbb{R}^{P\times M}$ has rank $d_{Y}+1$. The corresponding optimal weight matrices for Eq. (II.1) can be solved as $W_{*}=H_{*}X^{\dagger}$ (since $X$ is of full rank) and $W_{2,*}=YH_{*}^{T}[\lambda_{2}+H_{*}H_{*}^{T}]^{-1}$ (SI §1.2). So these matrices also have rank $d_{Y}+1$, which could be much smaller than the hidden layer width $M$ and input dimensions $d_{X}$. The fact that the optimal weight matrices are low-rank indicates that the neural network learns the task-relevant low-dimensional structures in the input, rather than constructs a large set of fixed features from the inputs as in kernel machine [97, 23].

### v2

@@ II.4 The optimality condition
To compute the $PS$ [Eq. (7)] corresponding to the ground state of effective system, we want to find the representation kernel $K[\rho_{M}^{*}]$ corresponding to the global minima $\rho_{M}^{*}$ of Eq. (14), with the constraint that $\rho_{M}^{*}$ is a sum of $M$ Dirac-delta measures and $\text{supp}\rho_{M}^{*}\subseteq\text{Range}K_{X}$. This optimization is challenging to perform directly, as the space of finite $M$-sum Dirac delta measures is not convex. To address this issue, we relax the constraint by enlarging the optimization domain to include all finite positive measures supported on $\text{Range}K_{X}\subseteq\mathbb{R}^{P}$, which also includes continuous measures. We denote this new space of measures  [^2] To be technically correct, we also require all 2nd-order moments of $\rho$ are finite in $M_{+}(X)$. This would ensure all the integrals in $E[\rho]$ are finite., as $M_{+}(K_{X})$. By definition, $M_{+}(K_{X})$ is convex (equipped with the usual sum between measures).
@@
This is a convex functional on $M_{+}(X)$: $E[\rho]$ is a convex function on $M_{+}(X)$ if and only if $\forall\rho_{1},\rho_{2}\in M_{+}(X),\,f(t)\equiv E[t\rho_{1}+(1-t)\rho_{2}]$ is a convex function for $t\in[0,1]$ [11]. The latter can be checked by computing the 2nd derivative of $f(t)$ (SI §1.3),
@@
$\mathbf{\,h}\in\text{supp}(\rho_{*})\,$ $\Rightarrow\,\text{equality holds,}$ (17)
where $K[\rho_{*}]=\int\phi(\mathbf{h})\phi(\mathbf{h})^{T}\,d\rho_{*}(\mathbf{h})$ is the representation kernel for $\rho_{*}$. Because the problem is convex in $\rho_{*}$, any solution to the KKT conditions is automatically a global minimum for Eq. (15).
This KKT condition shows that any global minimum $\rho_{*}$ of Eq. (15) must satisfy the inequality in Eq. (II.4) and at the same time, its support must be confined to those vectors $\mathbf{h}$ for which the equal sign is attained. The two conditions can be interpreted as a mean-field description [66] of the effective $M$-neuron system (Fig. 3F). To see this, we define the single-neuron mean-field energy,
@@
where $K[\rho]$ is the representation kernel for $\rho$, as defined above.
The KKT conditions [Eq. (II.4)] for any optimal solution $\rho_{*}$ can be written in terms of the single-neuron mean-field energy
@@
$\text{supp}(\rho_{*})\subseteq\underset{\mathbf{h}\in\text{Range}K_{X}}{\text{argmin}}E(\mathbf{h};\rho_{*})$
$\quad\quad\quad\quad\equiv A(K_{X},K_{Y}),$ (19)
In the second equation, the support of $\rho_{*}$ consists of the optimal preactivation patterns (or single-neuron tuning) for training data, which we denote as $A(K_{X},K_{Y})$.
@@
These mean-field equations [Eq. (II.4)] need to be solved in a self-consistent manner [66, 45, 46, 62] (See SI §1.4). For arbitrary input and output data $K_{X},K_{Y}$, and nonlinear activation function $\phi$, these equations are systems of nonlinear equations that can only be solved numerically. Fortunately, for training data related to abstract representation (Section II.1), the mean-field equations [Eq. (II.4)] can be solved exactly as we will present below.
As a final note, although the above optimization is performed over all positive measures in $M_{+}(X)$ and in general, some of these measures cannot be attained by a finite-size system, in many cases, the global minimizer $\rho_{*}$ is a finite sum of Dirac-delta measures. As we show in the next section, this fact allows us to study the optimal neural representations in a finite-width network by optimizing a convex functional [Eq. (15)] over the infinite-dimensional measure space for preactivation, $M_{+}(X)$.
@@ III Whitened and target-aligned inputs lead to abstract representation in ReLU network
For these input geometries, we find that all the solutions satisfying the KKT conditions [Eq. (II.4)] correspond to abstract hidden representations. We summarize the key steps of the argument here, with detailed derivations provided in SI §2. We start with the scenario where each binary class contains a single data point $(n=1)$ and then extend the results to multi-element classes $(n\geq 2)$.
@@ III.1 Whitened input + single-element class ($n=1$)
Both matrices are full, so the constraint on $\mathbf{h}$ in Eq. (II.4) is trivial, $\text{Range}K_{X}=\mathbb{R}^{P}$.
@@
where $\mathbf{h}_{+}=[\mathbf{h}]_{+}$ is the positive component of the vector $\mathbf{h}$. Using this inequality, we can find the minimum of $E(\mathbf{h};\rho)$ [as required in Eq.(II.4)] via minimizing $E_{r}(\mathbf{h}_{+};\rho)$ over all the nonnegative vectors $\mathbf{h}_{+}\geq 0$.
@@
Here $\lambda_{1,2}$ are assumed to be small to ensure $b_{*}>0$, i.e. $\lambda_{1}\lambda_{2}<\frac{P+2}{P(P+1)}$. Interestingly, the solutions of the mean-field equation $\rho_{*}$ [Eq.(II.4)], or equivalently the global minimizers of the loss [Eq.(15)], are not unique. But they all correspond to the identical representation kernel given by Eq.(21). We note that such uniqueness of the representation kernel does not generally hold for arbitrary tasks and appears to be a special property of the tasks studied here (SI §2.2)
Figure 4: Task-optimized ReLU network exhibits abstract representation for whitened and target-aligned inputs. (A) The training loss and the parallelism score of the hidden representation are plotted against the number of training epochs for the whitened input. The training is through gradient descent algorithm. After training, the network performs the task perfectly with zero training loss and has an abstract hidden representation ($PS\rightarrow 1$). (B) The optimal hidden kernel predicted by theory ($K_{theory}$ given by Eq. (21)) is aligned with the one found in numerical simulation ($K_{sim}$) for different output dimensions $d_{Y}$. (C) The parallelism score of the hidden representation after training vs. the input-output alignment. See SI §5 for the definition of input-output alignment. Each point in the plot represents a specific randomly sampled input geometry. The red dot indicates the point for the whitened input. For inputs that are more aligned to the output than the whitened one, the $PS$ is close to $1$. (D-E) Modularity of the single-neuron tuning in the hidden layer is captured by the preactivation vector $\mathbf{h}$ and weight vector $\mathbf{v}$ to the output layer for each hidden neuron ($d_{Y}=2$). (E) Each hidden neuron only has nonzero output weights to a single output unit, suggesting that different neurons in the hidden layer ”encode” different output labels. (F) The first three principal components of the neural preactivation space (the same space as Fig. 3B). The preactivation vectors of the hidden neurons concentrate along $P=4$ distinct directions, as predicted by Eq. (23).
The set of optimal preactivation vectors in the hidden layer $A(K_{X},K_{Y})$, which is equal to the support of $\rho_{*}$ , is found to consist of $2d_{Y}$ line rays in the nonnegative orthorant $\mathbb{R}_{\geq 0}^{P}$ (SI §2.2)
@@
As half of the components in $\mathbf{v}_{i}$ are +1+1 and the other ones are 1-1, the above optimal preactivation vectors have exactly half of the components to be $2$ and the other ones zero.
@@
Altogether, when $M\geq 2d_{Y}$ and the input is whitened, we find the optimal hidden representation is Eq.(21) and that the neurons in the hidden layer cluster into $2d_{Y}$ groups given by Eq.(23).
@@ III.2 Whitened input + multi-element class ($n\geq 2$)
In SI §2.3, we solve the mean-field problem [Eq. (II.4)] and find the optimal representation kernel is similar as before
@@
This optimal representation kernel [Eq. (25)] has an intriguing geometric interpretation: the between-class correlation matrix $b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y}^{0})$ is the same as for the single-element class case [Eq. (21)] and the within-class correlation matrix is given by $\mathbf{1}\mathbf{1}^{T}$. This form of within-class correlation means the neural representation of all data points from the same class “collapse” into a single point in the hidden layer, as also indicated in the set of optimal preactivations [Eq. (27)]. Such a property closely resembles the neural collapse phenomena reported in previous studies [68, 100, 92, 88, 87, 36]. We discuss the similarity and difference between our results and the work related to neural collapse in SI §5.1 in more detail. For ReLU nonlinearity, this ”within-class collapse” property can be directly proved via Jensen’s inequality (SI §2.4). The fact that the optimal preactivation patterns in the hidden layer clusters into only a few directions [Eq. (23) and Eq. (25)] is also known as the quantization phenomenon [54].
@@ III.3 Target-aligned input
Unlike the fully orthogonalized input before, the input here has positive (between-class) correlations if $c_{0}$ is large. Interestingly, we find (SI §2.6) that this introduced correlation does not change the form of the optimal kernel $K[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y})$, and only modifies the coefficient $b_{*}$
@@
The optimal kernel in this case has the same form as for orthogonalized inputs [Eq. (25)], $K[\rho_{*}]=b_{*}(d_{Y}\mathbf{1}\mathbf{1}^{T}+K_{Y}^{0})\otimes\mathbf{1}\mathbf{1}^{T}$, but with an coefficient (SI §2.6)
@@
Altogether, this shows that for target-aligned inputs [Eq.(29)], the optimal representation kernel [Eq. (28)] does not depend on the magnitude of orthogonal components $c_{j}$ for all $j=d_{Y}+1,..,P-1$. However, this property does not hold if the target-aligned condition [Eq.(29)] is not satisfied, i.e. when the orthogonal component is large, $c_{j}>c_{Y}$ (Fig.5C). The underlying intuition is that for nonlinear networks, the output-aligned component and the output-orthogonal component in the input ”compete” with each other to form the hidden representation. In contrast, a linear network only uses the output-aligned component to form the hidden representation.
@@ III.4 $PS$, single-neuron tuning and weight matrices for the optimal solution
From the optimal representation kernel, we can compute its $PS$, denoted $PS(K[\rho_{*}])$. By translation-and scale-invariance of $PS$ (Section II.2), $PS(K[\rho_{*}])=PS(K_{Y})$. By the definition of $K_{Y}$, the prototype representation [Eq. (6)] of each binary class corresponds to each vertex of a $d_{Y}$-dimensional hypercube (Fig. 2). So the representation change for the $k$th binary latent label is aligned with the $k$th axis of the hypercube (Fig.4), independent of other latent labels: $PS_{k}(\alpha_{1},\alpha_{2})\equiv\frac{\Delta\mathbf{r}(k;\alpha_{1})\cdot\Delta\mathbf{r}(k;\alpha_{2})}{\left\|\Delta\mathbf{r}(k;\alpha_{1})\right\|_{2}\left\|\Delta\mathbf{r}(k;\alpha_{2})\right\|_{2}}=1$. Therefore, the overall parallelism score $PS(K_{Y})=1=PS(K[\rho_{*}])$, and the optimal representation kernel [Eq. (24) and (31)] corresponds to an abstract representation.
Next, we look at the responses of individual neurons in the optimal solution $\rho_{*}$. From Eq. (23) and Eq. (27), the optimal preactivation of each neuron has the form $\mathbf{h}=\alpha(1\pm\mathbf{v}^{0}_{i})\otimes\mathbf{1}$ for some $i\in\{1,2,3..,d_{Y}\}$ and $\alpha>0$ (we focus on neurons having nonzero preactivation patterns). The $P$ components of $\mathbf{h}\in\mathbb{R}^{P}\simeq\mathbb{R}^{2^{d_{Y}}}\otimes\mathbb{R}^{n}$ are this neuron’s preactivations for the $P$ training data points [Eq.(11)]. Based on the definition of $\mathbf{v}^{0}_{i}$ (Section II.1 and SI §1.1), the neuron with $\mathbf{h}=\alpha(1+\mathbf{v}^{0}_{i})\otimes\mathbf{1}$ (or $\mathbf{h}=\alpha(1-\mathbf{v}^{0}_{i})\otimes\mathbf{1}$) has nonzero preactivations exactly for those training data points whose $i$th output label are positive (or negative). So all the neurons in the hidden layer are divided into $2d_{Y}$ groups: each group responds only to a single output label (Fig.5D-F). Note that a random network would not have a modular response property like this, and such modularity is a consequence of task training.
Since the preactivation of every neuron is a linear combination of $\mathbf{1\otimes\mathbf{1}}$ and $\mathbf{v}_{i}\otimes\mathbf{1}$ for some $i$, the optimal preactivation matrix $H_{*}$ [Eq. (11)] has the following form
@@
where $\mathbf{A}\in\mathbb{R}^{(d_{Y}+1)\times M}$ is the matrix encoding the coefficients $\alpha$ for all neurons. Importantly, $H_{*}\in\mathbb{R}^{P\times M}$ has rank $d_{Y}+1$. The corresponding optimal weight matrices for Eq. (II.1) can be solved as $W_{*}=H_{*}X^{\dagger}$ (since $X$ is of full rank) and $W_{2,*}=YH_{*}^{T}[\lambda_{2}+H_{*}H_{*}^{T}]^{-1}$ (SI §1.2). So these matrices also have rank $d_{Y}+1$, which could be much smaller than the hidden layer width $M$ and input dimensions $d_{X}$. The fact that the optimal weight matrices are low-rank indicates that the neural network learns the task-relevant low-dimensional structures in the input, rather than constructs a large set of fixed features from the inputs as in kernel machine [96, 22].